### v1

# Building a stable classifier with the inflated argmax

Abstract
We propose a new framework for algorithmic stability in the context of multiclass classification. In practice, classification algorithms often operate by first assigning a continuous score (for instance, an estimated probability) to each possible label, then taking the maximizer—i.e., selecting the class that has the highest score. A drawback of this type of approach is that it is inherently unstable, meaning that it is very sensitive to slight perturbations of the training data, since taking the maximizer is discontinuous. Motivated by this challenge, we propose a pipeline for constructing stable classifiers from data, using bagging (i.e., resampling and averaging) to produce stable continuous scores, and then using a stable relaxation of argmax, which we call the “inflated argmax”, to convert these scores to a set of candidate labels. The resulting stability guarantee places no distributional assumptions on the data, does not depend on the number of classes or dimensionality of the covariates, and holds for any base classifier. Using a common benchmark data set, we demonstrate that the inflated argmax provides necessary protection against unstable classifiers, without loss of accuracy.
1 Introduction
An algorithm that learns from data is considered to be stable if small perturbations of the training data do not lead to large changes in its output. Algorithmic stability is an important consideration in many statistical applications. Within the fairness literature, for instance, stability is one aspect of reliable decision-making systems (Friedler et al.,, 2019; Huang and Vishnoi,, 2019). In interpretable machine learning, it similarly serves as a form of reproducibility (Murdoch et al.,, 2019; Yu and Kumbier,, 2020; Yu and Barter,, 2024). Caramanis et al., (2011) relates stability to robustness, where a machine learning algorithm is robust if two samples with similar feature vectors have similar test error. In the context of generalization bounds, Bousquet and Elisseeff, (2002) and subsequent authors study stability of an algorithm’s real-valued output—for instance, values of a regression function. In the setting of a multiclass classification problem, where the data consists of features $X\in\mathcal{X}$ and labels $Y\in[L]=\{1,\dots,L\}$, the results of this literature can thus be applied to analyzing a classifier’s predicted probabilities $\hat{p}_{\ell}(X)\in[0,1]$—i.e., our learned estimates of the conditional probabilities $\Pr\{Y=\ell\mid X\}$, for each $\ell\in[L]$. However, as we will see, stability of these predicted probabilities by no means implies stability of the predicted label itself, $\hat{y}=\operatorname{argmax}_{\ell\in[L]}\hat{p}_{\ell}(x)$—an arbitrarily small perturbation in $\hat{p}_{\ell}(x)$ might completely change the predicted label $\hat{y}$. The distinction matters: for system trustworthiness, we care about the model’s final output, on which users base their decisions.
In this paper, we propose a new framework for algorithmic stability in the context of multiclass classification, to define—and achieve—a meaningful notion of stability when the output of the algorithm consists of predicted labels, rather than predicted probabilities. Our work connects to other approaches to allowing for ambiguous classification, including set-valued classification (Grycko,, 1993; Del Coz et al.,, 2009; Sadinle et al.,, 2019; Chzhen et al.,, 2021) and conformal inference (Papadopoulos et al.,, 2002; Vovk et al.,, 2005; Lei,, 2014; Angelopoulos et al.,, 2023) (we will discuss related work further, below).
1.1 Problem setting
In supervised classification, we take a data set $\mathcal{D}=\big{(}(X_{i},Y_{i})\big{)}_{i\in[n]}$ consisting of $n$ observations, and train a classifier that maps from the feature space $\mathcal{X}$ to the set $[L]$ of possible labels. [^1] We remark that all the results of this paper apply also to the case of countably infinitely many labels, $L=\infty$, in which case we should take $\mathbb{N}$ to be the label space instead of $[L]$. Typically, this map is constructed in two stages. First we run some form of regression to learn a map $\hat{p}:\mathcal{X}\to\Delta_{L-1}$ from features to predicted probabilities, with $\hat{p}_{\ell}(x)$ denoting our learned estimate of the conditional label probability, $\Pr\{Y=\ell\mid X=x\}$ (here $\Delta_{L-1}=\{w\in\mathbb{R}^{L}:w_{i}\geq 0,\sum_{i}w_{i}=1\}$ denotes the probability simplex in $L$ dimensions). We will write $\hat{p}=\mathcal{A}(\mathcal{D})$, where $\mathcal{A}$ denotes the learning algorithm mapping a data set $\mathcal{D}$ (of any size) to the map $\hat{p}$. Then the second step is to convert the predicted probabilities $\hat{p}_{\ell}(x)$ to a predicted label, which is typically done by taking the argmax, i.e., $\hat{y}=\operatorname{argmax}_{\ell}\hat{p}_{\ell}(x)$ (with some mechanism for breaking ties). This two-stage procedure can be represented in the following diagram:
Training data $\mathcal{D}=\big{(}(X_{i},Y_{i})\big{)}_{i\in[n]}$ Predicted probabilities $\hat{p}(x)\in\Delta_{L-1}$ Test features $x\in\mathcal{X}$ Predicted label $\hat{y}\in[L]$
Learning
algorithm $\mathcal{A}$
Selection
(via $\operatorname{argmax}$)
When predictions are ambiguous, meaning two or more classes nearly achieve the maximum predicted probability, the selected label becomes unstable and can change based on seemingly inconsequential perturbations to the training data. In other words, the above workflow is fundamentally incompatible with the goal of algorithmic stability. Consequently, in this paper we instead work with set-valued classifiers, which return a set of candidate labels—in practice, this typically leads to a singleton set for examples where we have high confidence in the label, but allows for a larger set in the case of ambiguous examples. While the idea of returning a set of candidate labels is not itself new, we will see that the novelty of our work lies in finding a construction that offers provable stability guarantees.
In this framework, a feature vector $x\in\mathcal{X}$ is now mapped to a set of candidate labels $\hat{S}\subseteq[L]$ (rather than a single label $\hat{y}$), as illustrated here:
Training data $\mathcal{D}=\big{(}(X_{i},Y_{i})\big{)}_{i\in[n]}$ Predicted probabilities $\hat{p}(x)\in\Delta_{L-1}$ Test features $x\in\mathcal{X}$ Candidate labels $\hat{S}\subseteq[L]$
Learning
algorithm $\mathcal{A}$
Selection
rule $s$
Formally, given a test point $x\in\mathcal{X}$, this two-stage procedure returns $\hat{S}=s(\hat{p}(x))\subseteq[L]$, where $\hat{p}=\mathcal{A}(\mathcal{D})$ is the output of the regression algorithm $\mathcal{A}$ trained on data $\mathcal{D}$. Here $s:\Delta_{L-1}\to\raisebox{1.79993pt}{\Large$\wp$}([L])$ denotes a selection rule, mapping a vector of estimated probabilities to the set of candidate labels. Of course, the earlier setting—where the procedure returns a single label $\hat{y}=\operatorname{argmax}_{\ell}\,\hat{p}_{\ell}(x)$, rather than a subset—can be viewed as a special case by simply taking $s$ to be the $\operatorname{argmax}$ operator.
If we instead allow for $\hat{S}$ to contain multiple candidate labels, a commonly used version of this framework is given by a top-5 procedure (or top-$k$, for any fixed $k$). That is, after running a learning algorithm $\mathcal{A}$ to estimate probabilities $\hat{p}(x)$, the selection rule $s$ then returns the top 5 labels—the labels $\ell_{1},\dots,\ell_{5}\in[L]$ corresponding to the highest 5 values of $\hat{p}_{\ell}(x)$. This approach is more stable than a standard argmax. However, the choice of 5 in this setting is somewhat arbitrary, and the set $\hat{S}$ always contains 5 candidate labels, regardless of the difficulty of the test sample—while intuitively, we might expect that it should be possible to return a smaller $\hat{S}$ for test points $x$ that are easier to classify (and a larger $\hat{S}$ if $x$ is more ambiguous). In contrast, in our work we seek a more principled approach, where we provide provable stability guarantees while also aiming for the set $\hat{S}$ to be as small as possible.
1.2 Contributions
In this work, we introduce new definitions of algorithmic stability in the context of classification, focusing on the stability of the predicted label. We reduce the problem of stabilizing a classifier to separately stabilizing the learning and selection stages, described above. For the selection step, we propose the inflated argmax operation, derive its stabilizing properties and give an algorithm to compute it efficiently. We combine this operation with bagging at the learning step, to provably stabilize any classifier. In particular, our guarantee holds with no assumptions on the data, and no constraints on the dimensionality of the covariates nor on the number of classes.
2 Framework: stable classification
In this section, we propose a definition of algorithmic stability in the setting of multiclass classification. To begin, we formally define a classification algorithm as a map [^2] We will assume without comment that $\mathcal{C}$ is measurable. Furthermore, in practice classification algorithms often incorporate randomization (either in the learning stage, such as via stochastic gradient descent, and/or in the selection stage, such as by using a random tie-breaking rule for $\operatorname{argmax}$). All of our definitions and results in this paper can be naturally extended to randomized algorithms—see Appendix B.
$\mathcal{C}:\cup_{n\geq 0}(\mathcal{X}\times[L])^{n}\ \times\ \mathcal{X}\ \longrightarrow\ \raisebox{1.79993pt}{\Large$\wp$}([L]),$
which maps a training data set $\mathcal{D}$ of any size $n$, together with a test feature vector $x\in\mathcal{X}$, to a candidate set of labels $\hat{S}=\mathcal{C}(\mathcal{D},x)$. To relate this notation to our earlier terminology, the two-stage selection procedure described in Section 1.1 corresponds to the classification algorithm
$\mathcal{C}(\mathcal{D},x)=s(\hat{p}(x))\textnormal{ where }\hat{p}=\mathcal{A}(\mathcal{D}).$
Abusing notation, we will write this as $\mathcal{C}=s\circ\mathcal{A}$, indicating that $\mathcal{C}$ is obtained by applying selection rule $s$ to the output of the learning algorithm $\mathcal{A}$.
We now present our definition of algorithmic stability for a classification algorithm $\mathcal{C}$. As is common in the algorithmic stability literature, we focus on stability of the algorithm’s output with respect to randomly dropping a data point: does the output of $\mathcal{C}$ on a test point $x$ change substantially if we drop a single data point from the training data $\mathcal{D}$? If the output of the algorithm were a real-valued prediction $\hat{y}\in\mathbb{R}$, we could assess this by measuring the real-valued change in $\hat{y}$ when a single data point is dropped. For classification, however, we will need to take a different approach:
Definition 1.
We say a classification algorithm $\mathcal{C}$ has selection stability $\delta$ at sample size $n$ if, for all datasets $\mathcal{D}\in(\mathcal{X}\times[L])^{n}$ and all test features $x\in\mathcal{X}$,
$\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\left\{\hat{S}\cap\hat{S}^{\setminus i}=\varnothing\right\}\leq\delta,$
where $\hat{S}=\mathcal{C}(\mathcal{D},x)$ and where $\hat{S}^{\setminus i}=\mathcal{C}(\mathcal{D}^{\setminus i},x)$, for each $i\in[n]$.
Here the notation $\mathcal{D}^{\setminus i}$ denotes the data set $\mathcal{D}$ with $i$th data point removed—that is, for a data set $\mathcal{D}=\big{(}(X_{j},Y_{j})\big{)}_{j\in[n]}$, the leave-one-out data set is given by $\mathcal{D}^{\setminus i}=\big{(}(X_{j},Y_{j})\big{)}_{j\in[n]\setminus i}$.
2.1 Interpreting the definition
Intuitively, selection stability controls the probability that our algorithm makes wholly inconsistent claims when dropping a single data point at random from the training set. If we interpret $\hat{S}$ as making the claim that the true label $Y$ is equal to one of the labels in the candidate set $\hat{S}$, and similarly for $\hat{S}^{\setminus i}$, then as soon as there is even one single value that lies in both $\hat{S}$ and $\hat{S}^{\setminus i}$, this means that the two claims are not contradictory—even if the sets are large and are mostly nonoverlapping.
At first glance, this condition appears to be too mild, in the sense that we require the two prediction sets only to have some way of being compatible, and allows for substantial difference between the sets $\hat{S}$ and $\hat{S}^{\setminus i}$. However, since standard classification algorithms always output a single label, they often cannot be said to be stable even in this basic sense. Thus, we can view this definition as providing a minimal notion of stability that we should require any interpretable method to satisfy.
2.2 Connection to classical algorithmic stability
Most prior work on algorithmic stability concerns algorithms with continuous outputs—for example, in our notation above, the algorithm $\mathcal{A}$ that returns estimated probabilities $\hat{p}(x)$. With such algorithms, there are more standard tools at our disposal for quantifying and ensuring stability. In this section, we connect classical algorithmic stability to our notion of selection stability (Definition 1). We first recall the following definition.
Definition 2.
(Soloff et al., 2024b, ) A learning algorithm $\mathcal{A}$ has tail stability $(\varepsilon,\delta)$ at sample size $n$ if, for all datasets $\mathcal{D}$ of size $n$ and all test features $x\in\mathcal{X}$,
$\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}\left\{\|\hat{p}(x)-\hat{p}^{\setminus i}(x)\|\geq\varepsilon\right\}\leq\delta,$
[…]
This intuitive notion of stability is achieved by many well known algorithms $\mathcal{A}$, such as nearest-neighbor type methods, or methods based on bagging or ensembling (as established by Soloff et al., 2024a ; Soloff et al., 2024b —see Section 3.1 below for details).
2.3 From stability to selection stability
To construct a classification algorithm using the two-stage procedure outlined in Section 1.1, we need to apply a selection rule $s$ to the output of our learning algorithm $\mathcal{A}$. We might expect that choosing a stable $\mathcal{A}$ will lead to selection stability in the resulting classification algorithm—but in fact, this is not the case: even if the learning algorithm $\mathcal{A}$ is itself extremely stable in the sense of Definition 2, the classification rule obtained by applying $\operatorname{argmax}$ to the output of $\mathcal{A}$ can still be extremely unstable. The underlying issue is that $\operatorname{argmax}$ is extremely discontinuous—the perturbation $\|\hat{p}(x)-\hat{p}^{\setminus i}(x)\|$ in the predicted probabilities can be arbitrarily small but still lead to different predicted labels, i.e., $\operatorname{argmax}_{\ell}\hat{p}_{\ell}(x)\neq\operatorname{argmax}_{\ell}\hat{p}^{\setminus i}_{\ell}(x)$.
Since combining $\operatorname{argmax}$ with a stable learning algorithm $\mathcal{A}$ will not suffice, we instead seek a different selection rule $s$—one that will ensure selection stability (when paired with a stable learning algorithm $\mathcal{A}$). To formalize this aim, we introduce another definition:
Definition 3.
A selection rule $s:\Delta_{L-1}\to\raisebox{1.79993pt}{\Large$\wp$}([L])$ is $\varepsilon$-compatible if, for any $v,w\in\Delta_{L-1}$,
$\|v-w\|<\varepsilon\ \Longrightarrow\ s(v)\cap s(w)\neq\varnothing.$
This notion of $\varepsilon$-compatibility allows us to construct classification algorithms with selection stability. Combining the above definitions leads immediately to the following result:
Proposition 4.
Let $\mathcal{A}$ be a learning algorithm with tail stability $(\varepsilon,\delta)$ at sample size $n$, and let $s$ be a selection rule satisfying $\varepsilon$-compatibility. Then the classification algorithm $\mathcal{C}=s\circ\mathcal{A}$ has selection stability $\delta$ at sample size $n$.
Therefore, by pairing a stable learning algorithm $\mathcal{A}$ with a compatible selection rule $s$, we will automatically ensure selection stability of the resulting classification algorithm $\mathcal{C}=s\circ\mathcal{A}$.
Of course, $\varepsilon$-compatibility of the selection rule $s$ might be achieved in a trivial way—for instance, $s$ returns the entire set $[L]$ for any input. As is common in statistical settings (e.g., a tradeoff between Type I error and power, in hypothesis testing problems), our goal is to ensure selection stability while returning an output $\hat{S}$ that is as informative as possible. In particular, later on we will consider the specific aim of constructing $\hat{S}$ to be a singleton set as often as possible.
3 Methodology: assumption-free stable classification
In this section, we formally define our pipeline for building a stable classification procedure using any base learning algorithm $\mathcal{A}$. At a high level, we leverage 4 and separately address the learning and selection stages described in Section 2.
  1. 1.
    In Section 3.1, we construct a bagged (i.e., ensembled) version of the base learning algorithm $\mathcal{A}$. The recent work of Soloff et al., 2024b ensures that the resulting bagged algorithm has tail stability $(\varepsilon,\delta)$, with $\varepsilon\asymp\frac{1}{\sqrt{n\delta}}$.
  2. 2.
    In Section 3.2, we introduce a new selection rule, the inflated argmax, and establish its $\varepsilon$-compatibility. Combined with the bagged algorithm, then, the resulting classification algorithm will be guaranteed to have selection stability $\delta$ (by 4).
Before describing these two steps in detail, we first present our main theorem that characterizes the selection stability guarantee of the resulting procedure. Given sample size $n$ for the training data set $\mathcal{D}$, the notation $\widetilde{A}_{m}$ will denote a bagged version of the base algorithm $\mathcal{A}$, obtained by averaging over subsamples of $\mathcal{D}$ comprised of $m$ data points sampled either with replacement (“bootstrapping”, commonly run with $m=n$) or without replacement (“subbagging”, commonly run with $m=n/2$)—see below for details.
Theorem 5.
Fix any sample size $n$, any bag size $m$, and any inflation parameter $\varepsilon>0$. For any base learning algorithm $\mathcal{A}$, the classification algorithm $\mathcal{C}=\operatorname{argmax}^{\varepsilon}\circ\widetilde{A}_{m}$, obtained by combining the bagged version of $\mathcal{A}$ together with the inflated argmax, satisfies selection stability $\delta$ where
$\delta=\varepsilon^{-2}\cdot\frac{1-1/L}{n-1}\cdot\frac{p_{n,m}}{1-p_{n,m}},$ (1)
where $p_{n,m}=1-(1-\frac{1}{n})^{m}$ for bootstrapping, and $p_{n,m}=\frac{m}{n}$ for subbagging.
The guarantee in Theorem 5 holds for any base learning algorithm $\mathcal{A}$, and applies regardless of the complexity of the feature space $\mathcal{X}$, and allows the test feature $x\in\mathcal{X}$ to be chosen adversarially. The dependence on the number of classes $L$ is mild—in fact, the tail stability parameter $\delta$ in (1) differs only by a factor of two for $L=2$ versus $L=\infty$. Of course, the guarantee does depend on the choice of the bag size $m$. In general a smaller value of $m$ leads to a stronger stability guarantee (since $p_{n,m}$ increases with $m$), but this comes at the expense of accuracy since we are training the base algorithm $\mathcal{A}$ on subsampled data sets $\mathcal{D}^{r}$ of size $m$ (rather than $n$). For the common choices of $m=n$ for bootstrap or $m=n/2$ for subbagging, we have $\frac{p_{n,m}}{1-p_{n,m}}=\mathcal{O}(1)$ for each.
3.1 Bagging classifier weights
In this section, we formally define the construction of a bagged algorithm. We consider the two most common variants of this ensembling method: bagging (based on bootstrapping training points) and subbagging (based on subsampling). For any data set $\mathcal{D}\in(\mathcal{X}\times[L])^{n}$, we can define a subsampled data set of size $m$ as follows: for a sequence $r=(i_{1},\dots,i_{m})\in[n]^{m}$ (which is called a bag), we define the corresponding subsampled data set $\mathcal{D}^{r}=\big{(}(X_{i},Y_{i})\big{)}_{i\in r}\in(\mathcal{X}\times[L])^{m}$. Note that if the bag $r$ contains repeated indices (i.e., $i_{k}=i_{\ell}$ for some $k\neq\ell$), then the same data point from the original data set $\mathcal{D}$ will appear multiple times in $\mathcal{D}^{r}$.
Definition 6 (Bootstrapping or subbagging a base learning algorithm $\mathcal{A}$).
Given a data set $\mathcal{D}\in(\mathcal{X}\times[L])^{n}$ and some $x\in\mathcal{X}$, return the output $\widetilde{\mathcal{A}}_{m}(\mathcal{D})(x)=\mathbb{E}_{r}[\mathcal{A}(\mathcal{D}^{r})(x)]$, where the expected value is taken with respect to a bag bag $r$ that is sampled as follows:
  • Bootstrapping (sometimes simply referred to as bagging) (Breiman, 1996a, ; Breiman, 1996b, ) constructs each bag $r$ by sampling $m$ indices $r=(i_{1},\dots,i_{m})$ uniformly with replacement from $[n]$.
  • Subbagging (Andonova et al.,, 2002) constructs each bag $r$ by sampling $m\leq n$ indices $r=(i_{1},\dots,i_{m})$ uniformly without replacement from $[n]$.
The following result (Soloff et al., 2024b, ) ensures tail stability for any bootstrapped or subbagged algorithm:
Theorem 7 ((Soloff et al., 2024b, )).
For any base learning algorithm $\mathcal{A}$ returning outputs in $\Delta_{L-1}$, the bagged algorithm $\widetilde{\mathcal{A}}_{m}$ has tail stability $(\varepsilon,\delta)$ for any $\varepsilon,\delta>0$ satisfying (1).
Computational cost of bagging.
In this section, we have worked with the ideal, derandomized version of bagging for simplicity—that is, we assume that the expected value $\mathbb{E}_{r}[\mathcal{A}(\mathcal{D}^{r})(x)]$ is calculated exactly with respect to the distribution of $r$. In practice, of course, this is computationally infeasible (since for bootstrapping, say, there are $n^{m}$ possible bags $r$), and so we typically resort to Monte Carlo sampling to approximate this expected value, defining $\widetilde{\mathcal{A}}_{m}(\mathcal{D})(x)=\frac{1}{B}\sum_{b=1}^{B}\mathcal{A}(\mathcal{D}^{r_{b}})(x)$, for $B$ i.i.d. bags $r_{1},\dots,r_{B}$ sampled via bootstrapping or subbagging. See Appendix C for extensions of our main result, Theorem 5, to the finite-$B$ version of the method.
3.2 The inflated argmax
We are now ready to define the inflated argmax, denoted by $\operatorname{argmax}^{\varepsilon}$, which will be used as our selection rule $s$ for building our classification algorithm. First, for each $j\in[L]$, define
$R_{j}^{\varepsilon}=\left\{w\in\mathbb{R}^{L}:w_{j}\geq\max_{\ell\neq j}w_{\ell}+\varepsilon/\sqrt{2}\right\},$
which is a closed and convex subset of $\mathbb{R}^{L}$. For any $w\in\mathbb{R}^{L}$, we then define
$\operatorname{argmax}^{\varepsilon}(w):=\left\{j\in[L]:\textnormal{dist}(w,R_{j}^{\varepsilon})<\varepsilon\right\},$ (2)
where $\textnormal{dist}(w,R_{j}^{\varepsilon})=\inf_{v\in R_{j}^{\varepsilon}}\|w-v\|$. Intuitively, $j\in\operatorname{argmax}^{\varepsilon}(w)$ indicates that a small perturbation of $w$ would lead to the $j$th entry being largest by some margin.
While $\operatorname{argmax}^{\varepsilon}$ is defined as an operator on $\mathbb{R}^{L}$, in the context of classification algorithms, we only apply it to vectors $w$ lying in the simplex $\Delta_{L-1}$ (in particular, in the context of a two-stage classification procedure as in Section 1.1, we will apply the inflated argmax to the vector $w=\hat{p}(x)$ of estimated probabilities). In Figure 1, we visualize the inflated argmax applied to the simplex in a setting with three possible labels, $L=3$. Each different shaded area corresponds to the region of the simplex $\Delta_{L-1}$ where $\operatorname{argmax}^{\varepsilon}(w)$ returns a particular subset.
Refer to captionRefer to caption
Figure 1: The left plot illustrates the inflated argmax (2) over the simplex $\Delta_{L-1}$ when $L=3$. The numbers in brackets corresponds to the output of the inflated argmax, $\operatorname{argmax}^{\varepsilon}(w)$, for various points $w$ in the simplex. The right plot shows the same but for the standard argmax, which corresponds to the limit of $\operatorname{argmax}^{\varepsilon}(w)$ as $\varepsilon\rightarrow 0$.
3.2.1 Inflated argmax leads to selection stability
Unlike the standard argmax, the inflated argmax allows us to achieve stability in the context of classification. This is due to the following theorem, which shows that the inflated argmax is $\varepsilon$-compatible, as introduced in Definition 3. (The proofs of this result and all properties of the inflated argmax are deferred to Appendix A.)
Theorem 8.
For any $\varepsilon>0$ and any $v,w\in\mathbb{R}^{L}$, if $\|v-w\|<\varepsilon$ then $\operatorname{argmax}^{\varepsilon}(v)\cap\operatorname{argmax}^{\varepsilon}(w)\neq\varnothing$. In particular, viewing $\operatorname{argmax}^{\varepsilon}$ as a map on the simplex $\Delta_{L-1}$, $\operatorname{argmax}^{\varepsilon}$ is $\varepsilon$-compatible.
To gain a more concrete understanding of this theorem, we can look again at Figure 1 to examine the case $L=3$. Theorem 8 ensures that any two regions corresponding to disjoint subsets—e.g., the top corner region corresonding to $\{2\}$, and the bottom center region corresponding to $\{1,3\}$—are distance at least $\varepsilon$ apart. In particular, the region in the center, corresponding to vectors $w$ that map to the full set $\{1,2,3\}$, is a curved triangle of constant width $\varepsilon$, also known as a Reuleaux triangle (Reuleaux,, 1876; Weisstein,, 2005).
Proof of Theorem 5.
With the above results in place, we can now see that the inflated argmax allows us to achieve selection stability, when combined with a bagged algorithm. In particular, combining Theorem 7, Theorem 8, and 4 immediately implies our main result, Theorem 5.
3.2.2 Additional properties of the inflated argmax
The following result establishes some natural properties obeyed by the inflated argmax, and also compares to the standard argmax.
Proposition 9 (Basic properties of the inflated argmax).
Fix any $\varepsilon>0$. The inflated argmax operator satisfies the following properties:
  1. 1.
    (Including the argmax.) For any $w\in\mathbb{R}^{L}$, $\operatorname{argmax}(w)\subseteq\operatorname{argmax}^{\varepsilon}(w)$. [^3] In this result, $\operatorname{argmax}(w)$ should be interpreted as the set of all maximizing entries of $w$—i.e., in the case of ties, $\operatorname{argmax}(w)$ may not be a singleton set. Moreover, $\cap_{\varepsilon>0}\operatorname{argmax}^{\varepsilon}(w)=\operatorname{argmax}(w)$.
  2. 2.
    (Monotonicity in $\varepsilon$.) For any $w\in\mathbb{R}^{L}$ and any $\varepsilon<\varepsilon^{\prime}$, $\operatorname{argmax}^{\varepsilon}(w)\subseteq\operatorname{argmax}^{\varepsilon^{\prime}}(w)$.
  3. 3.
    (Monotonicity in $w$.) For any $w\in\mathbb{R}^{L}$, if $w_{j}\leq w_{k}$, then $j\in\operatorname{argmax}^{\varepsilon}(w)$Rightarrow $k\in\operatorname{argmax}^{\varepsilon}(w)$.
  4. 4.
    (Permutation invariance.) For any $v,w\in\mathbb{R}^{L}$, if $v=(w_{\sigma(1)},\dots,w_{\sigma(L)})$ for some permutation $\sigma$ on $[L]$, then $j\in\operatorname{argmax}^{\varepsilon}(v)$Leftrightarrow $\sigma(j)\in\operatorname{argmax}^{\varepsilon}(w)$.
Next, while we have established that inflated argmax offers favorable stability properties, we have not yet asked whether it can be efficiently computed—in particular, it is not immediately clear how to verify the condition $\textnormal{dist}(w,R_{j}^{\varepsilon})<\varepsilon$ in (2), in order to determine whether $j\in\operatorname{argmax}^{\varepsilon}(w)$. The following result offers an efficient algorithm for computing the inflated argmax set.
Proposition 10 (Computing the inflated argmax).
Fix any $w\in\mathbb{R}^{L}$ and $\varepsilon>0$. Let $w_{[1]}\geq\dots\geq w_{[L]}$ denote the order statistics of $w$, and define
$\hat{k}(w)=\max\bigg{\{}k\in[L]:\Big{(}\sum_{\ell=1}^{k}(w_{[\ell]}-w_{[k]})\Big{)}^{2}+\sum_{\ell=1}^{k}(w_{[\ell]}-w_{[k]})^{2}\leq\varepsilon^{2}\bigg{\}}.$
Then
$\operatorname{argmax}^{\varepsilon}(w)=\left\{j\in[L]:w_{j}>\frac{\varepsilon}{\sqrt{2}}+\hat{A}_{1}(w)-\sqrt{\hat{k}(w)+1}\sqrt{\frac{\varepsilon^{2}}{\hat{k}(w)}+(\hat{A}_{1}(w))^{2}-\hat{A}_{2}(w)}\right\},$
where $\hat{A}_{1}(w)=\frac{w_{[1]}+\dots+w_{[\hat{k}(w)]}}{\hat{k}(w)}$, and $\hat{A}_{2}(w)=\frac{w_{[1]}^{2}+\dots+w_{[\hat{k}(w)]}^{2}}{\hat{k}(w)}$.
Finally, thus far we have established that $\operatorname{argmax}^{\varepsilon}$ enables us to achieve stable classification with a computationally efficient selection rule—but we do not yet know whether $\operatorname{argmax}^{\varepsilon}$ is optimal, or whether some other choice of $s$ might lead to smaller output sets $\hat{S}$ (while still offering assumption-free stability). For instance, we might consider a fixed-margin rule,
$s_{\textnormal{margin}}^{\varepsilon}(w)=\{j:w_{j}>\max_{\ell}w_{\ell}-\varepsilon/\sqrt{2}\},$ (3)
for which $\varepsilon$-compatibility also holds—might this simpler rule be better than the inflated argmax? Our next result establishes that—under some natural constraints—$\operatorname{argmax}^{\varepsilon}$ is in fact the optimal choice of selection rule, in the sense of returning a singleton set (i.e., $|\hat{S}|=1$) as often as possible.
Proposition 11 (Optimality of the inflated argmax).
Let $s:\Delta_{L-1}\to\raisebox{1.79993pt}{\Large$\wp$}([L])$ be any selection rule. Suppose $s$ is $\varepsilon$-compatible (Definition 3), permutation invariant (in the sense of 9), and contains the argmax. Then for any $w\in\Delta_{L-1}$ and any $j\in[L]$,
$s(w)=\{j\}\ \Longrightarrow\ \operatorname{argmax}^{\varepsilon}(w)=\{j\}.$
In other words, for any selection rule $s$ satisfying the assumptions of the proposition (which includes the fixed-margin rule, $s_{\textnormal{margin}}^{\varepsilon}$), if $s(w)$ is a singleton set then so is $\operatorname{argmax}^{\varepsilon}(w)$. (See also Appendix D for a closer comparison between the inflated argmax and the fixed-margin selection rule given in (3).)
4 Experiments
Figure 2: Results on the Fashion MNIST data set. The figure shows the instability $\delta_{j}$ (defined in (4)) over all test points $j=1,\dots,N$. The curves display the fraction of $\delta_{j}$’s that exceed $\delta$, for each value $\delta\in[0,1]$. The vertical axis is on a log scale. See Section 4 for details.
In this section, we evaluate our proposed pipeline, combining subbagging with the inflated argmax, with deep learning models and on a common benchmark data set. [^4] Code to fully reproduce the experiment is available at https://github.com/jake-soloff/stable-argmax-experiments. Training all of the models for this experiment took a total of four hours on 10 CPUs running in parallel on a single computing cluster.
Data and models.
We use Fashion-MNIST (Xiao et al.,, 2017), which consists of $n=60,000$ training pairs $(X_{i},Y_{i})$, $N=10,000$ test pairs $(\tilde{X}_{j},\tilde{Y}_{j})$, and $L=10$ classes. For each data point $(X,Y)$, $X$ is a $28\times 28$ grayscale image that pictures a clothing item, and $Y\in[L]$ indicates the type of item, e.g., a dress, a coat, etc. The base model we use is a variant of LeNet-5, implemented in PyTorch (Paszke et al.,, 2019) tutorials as GarmentClassifier(). The base algorithm $\mathcal{A}$ trains this classifier using $5$ epochs of stochastic gradient descent.
Methods and evaluation.
We compare three methods:
  1. 1.
    Taking the argmax of the base learning algorithm $\mathcal{A}$.
  2. 2.
    Taking the argmax of the subbagged algorithm $\widetilde{\mathcal{A}}_{m}$, with $B=200$ bags of size $m=n/2$.
  3. 3.
[…]
We evaluate each method based on several metrics. First, to assess selection stability, we compute, for each test point $j=1,\ldots,N$,
$\delta_{j}:=\frac{1}{500}\sum_{k=1}^{500}\mathbf{1}\left\{s(\hat{p}(\tilde{X}_{j}))\cap s(\hat{p}^{\setminus i_{k}}(\tilde{X}_{j}))\neq\varnothing\right\},$ (4)
where $i_{1},\dots,i_{500}$ are sampled uniformly without replacement from $[n]$ (i.e., we are sampling 500 leave-one-out models $\hat{p}^{\setminus i}$ to compare to the model $\hat{p}$ trained on the full training set). Second, to evaluate how accurate each method is, we compute average precision (how often the method unambiguously returns the correct label) and set size (the number of labels in the candidate set):
$\beta_{\textnormal{prec}}=\frac{1}{N}\sum_{j=1}^{N}\mathbf{1}\left\{s(\hat{p}(\tilde{X}_{j}))=\{\widetilde{Y}_{j}\}\right\},\quad\beta_{\textnormal{size}}=\frac{1}{N}\sum_{j=1}^{N}\big{|}s(\hat{p}(\tilde{X}_{j}))\big{|}.$
Ideally we would want a method to achieve high precision (a value of $\beta_{\textnormal{prec}}\in[0,1]$ that is close to $1$) and low set size (a value of $\beta_{\textnormal{size}}\geq 1$ that is close to $1$).
Results.
In Figure 2 and Table 1, we present results for the instability and accuracy of each method. In the figure, we show the survival function of $(\delta_{j})_{j\in[N]}$. The base algorithm, $\operatorname{argmax}\circ\mathcal{A}$, has the longest tail, i.e., $\delta_{j}$ is large for many test points $j$. The other two methods, $\operatorname{argmax}\circ\widetilde{\mathcal{A}}_{m}$ and $\operatorname{argmax}^{\varepsilon}\circ\widetilde{\mathcal{A}}_{m}$, both improve upon this tail, but the latter method shows the best stability by far: using the inflated argmax ensures that even the largest $\delta_{j}$ is small, with $\max_{j}\delta_{j}\approx 0.14$. In the table, we display the average accuracy measures, $\beta_{\textnormal{prec}}$ and $\beta_{\textnormal{size}}$. We see that precision is nearly identical for all three methods. Moreover, while the first two methods always return a set of size 1 (since we are using the standard argmax), the inflated argmax increases set size only very slightly in this experiment. In other words, the inflated argmax allows us to achieve substantially higher stability without sacrificing accuracy in this high signal data set.
$\beta_{\textnormal{prec}}$ $\beta_{\textnormal{size}}$
$\operatorname{argmax}\circ\mathcal{A}$ 0.879 (0.003) 1.000 (0.000)
$\operatorname{argmax}\circ\widetilde{\mathcal{A}}_{m}$ 0.893 (0.003) 1.000 (0.000)
$\operatorname{argmax}^{\varepsilon}\circ\widetilde{\mathcal{A}}_{m}$ 0.886 (0.003) 1.018 (0.001)
Table 1: Results on the Fashion MNIST data set. The table displays the average precision, $\beta_{\textnormal{prec}}$, and size, $\beta_{\textnormal{size}}$, as defined in Section 4. Standard errors are shown in parentheses.
5 Discussion
Related work.
There is an enormous literature on set-valued classification, so we only reference some of the most relevant works. Much prior work has considered the possibility of a more relaxed argmax to allow for uncertainty (Bridle,, 1990; de Brébisson and Vincent,, 2016; Martins and Astudillo,, 2016; Blondel et al.,, 2020; Zhao et al.,, 2023). The more recent literature has focused on producing a sparse set of weights, but none of these works offer a formal stability guarantee for the support of the weights. Our work is the first to propose a version of the argmax that can provably control a notion of stability for the classification setting.
Our definition of selection stability relies on a notion of consistency between sets—two sets of candidate labels are consistent if they have at least one common element. This is similar in flavor to conformal classification (Lei,, 2014; Sadinle et al.,, 2019), where a set of candidate values is valid if it contains the correct label; this method does not rely on any distributional assumptions, and consequently has been applied to complex and high-dimensional applications such as generative language models (Quach et al.,, 2023). These frameworks share the motivation of avoiding ‘overconfidence in the assignment of a definite label’ (Hechtlinger et al.,, 2018).
Overview, limitations and open questions.
We prove a guarantee on the selection stability based on our methodology combining bagging with the inflated argmax. Theorem 5 does not place any assumptions on the learning algorithm nor on the data, including any distributional assumptions. In fact, the training data and test point may be chosen adversarially based on the algorithm, and the output will still be stable. Moreover, we do not assume that the sample size is large: the guarantee holds for any fixed training set size $n$ and improves as $n$ increases. Furthermore, the inflated argmax selection rule ensures that the returned sets of candidate labels are as small as possible (i.e., that there is as little ambiguity as possible about the predicted label for any given test point $x$).
While our theorem does not require assumptions, our method does require bagging the base learning algorithm. The main limitation of our work is that bagging is computationally intensive. However, training different bags is easily parallelizable, which is what allowed us to easily train a convolutional neural network on $1,000$ total subsets of the training data.
We leave open several avenues for future work. Practitioners may be interested in other forms of discrete stability, which are more strict than requiring one common element between $\hat{S}$ and $\hat{S}^{\setminus i}$. For instance, one popular way to measure set similarity is the Jaccard index, given by $\frac{|\hat{S}\cap\hat{S}^{\setminus i}|}{|\hat{S}\cup\hat{S}^{\setminus i}|}$. Another open problem is to extend the stability framework and methods studied here to more structured discrete outputs, such as in classification.
Acknowledgements
We gratefully acknowledge the support of the National Science Foundation via grant DMS-2023109 DMS-2023109. R.F.B. and J.A.S. gratefully acknowledge the support of the Office of Naval Research via grant N00014-20-1-2337. R.M.W. gratefully acknowledges the support of NSF DMS-2235451, Simons Foundation MP-TMPS-00005320, and AFOSR FA9550-18-1-0166. The authors thank Melissa Adrian for help with running the experiments.
Appendix A Proofs of properties of the inflated argmax
In this section, we prove all the results of Section 3.2, establishing the various properties of the inflated argmax. In fact, all the proofs will rely on the following theorem, which gives an alternative characterization of the set $\operatorname{argmax}^{\varepsilon}(w)$.
Theorem 12.
Fix $\varepsilon>0$. Define the map $c_{\varepsilon}:\mathbb{R}^{L}\rightarrow\mathbb{R}$ as
$c_{\varepsilon}(w)\textnormal{ is the unique solution to }\bigg{(}\sum_{j\in[L]}(w_{j}-c)_{+}\bigg{)}^{2}+\sum_{j\in[L]}(w_{j}-c)^{2}_{+}=\varepsilon^{2},$
and define the map $t_{\varepsilon}:\mathbb{R}^{L}\rightarrow\mathbb{R}$ as
$t_{\varepsilon}(w)=c_{\varepsilon}(w)+\varepsilon/\sqrt{2}-\sum_{j\in[L]}(w_{j}-c_{\varepsilon}(w))_{+}.$
Then for any $w\in\mathbb{R}^{d}$, it holds that
$\operatorname{argmax}^{\varepsilon}(w)=\{j\in[L]:w_{j}>t_{\varepsilon}(w)\}.$
Moreover, $t_{\varepsilon}(w)\leq c_{\varepsilon}(w)$ for all $w$ and all $\varepsilon$, and $t_{\varepsilon}(w)$ and $c_{\varepsilon}(w)$ are both nondecreasing functions of $\varepsilon$, and both nondecreasing functions of $w$ (i.e., if $w\leq w^{\prime}$ holds coordinatewise, then $c_{\varepsilon}(w)\leq c_{\varepsilon}(w^{\prime})$ and $t_{\varepsilon}(w)\leq t_{\varepsilon}(w^{\prime})$).
With this key result in place, we are now ready to turn to the proofs of the results stated in Section 3.2. Throughout all the proofs below, the functions $c_{\varepsilon}(w)$ and $t_{\varepsilon}(w)$ are defined as in the statement of Theorem 12.
A.1 Proof of Theorem 8
Let $w,w^{\prime}\in\mathbb{R}^{L}$ be any vectors with $\operatorname{argmax}^{\varepsilon}(w)\cap\operatorname{argmax}^{\varepsilon}(w^{\prime})=\varnothing$. Define
$B=\{j\in[L]:w_{j}>c_{\varepsilon}(w)\},\quad B^{\prime}=\{j\in[L]:w^{\prime}_{j}>c_{\varepsilon}(w^{\prime})\},$
and note that we must have $B\subseteq\operatorname{argmax}^{\varepsilon}(w)$ since for $j\not\in\operatorname{argmax}^{\varepsilon}(w)$, $w_{j}\leq t_{\varepsilon}(w)\leq c_{\varepsilon}(w)$ by Theorem 12. Similarly, $B^{\prime}\subseteq\operatorname{argmax}^{\varepsilon}(w^{\prime})$. We also must have $B$ (and similarly $B^{\prime}$) nonempty by definition of $c_{\varepsilon}(w)$.
For convenience, define $c=c_{\varepsilon}(w)$, $c^{\prime}=c_{\varepsilon}(w^{\prime})$, $t=t_{\varepsilon}(w)$, $t^{\prime}=t_{\varepsilon}(w^{\prime})$. We then have
$\|w-w^{\prime}\|^{2}\geq\|w_{B}-w^{\prime}_{B}\|^{2}+||w^{\prime}_{B^{\prime}}-w_{B^{\prime}}\|^{2}\\ =\|(w_{B}-c\mathbf{1}_{B})+(t^{\prime}\mathbf{1}_{B}-w^{\prime}_{B})+(c-t^{\prime})\mathbf{1}_{B}\|^{2}+\|(w^{\prime}_{B^{\prime}}-c^{\prime}\mathbf{1}_{B^{\prime}})+(t\mathbf{1}_{B^{\prime}}-w_{B^{\prime}})+(c^{\prime}-t)\mathbf{1}_{B^{\prime}}\|^{2},$ (5)
where the first step holds since $B\cap B^{\prime}=\varnothing$ due to $\operatorname{argmax}^{\varepsilon}(w)\cap\operatorname{argmax}^{\varepsilon}(w^{\prime})=\varnothing$. Next we will need a technical lemma.
Lemma 13.
Let $k,k^{\prime}\geq 1$ be integers and let $\varepsilon>0$. Then for any $x,y\in\mathbb{R}^{k}_{\geq 0}$, $x^{\prime},y^{\prime}\in\mathbb{R}^{k^{\prime}}_{\geq 0}$, and $r,r^{\prime}\in\mathbb{R}$, if
$(x^{\top}\mathbf{1}_{k})^{2}+\|x\|^{2}=(x^{\prime}{}^{\top}\mathbf{1}_{k^{\prime}})^{2}+\|x^{\prime}\|^{2}=\varepsilon^{2}$ (6)
and
$x^{\top}\mathbf{1}_{k}+x^{\prime}{}^{\top}\mathbf{1}_{k^{\prime}}=r+r^{\prime}+\varepsilon\sqrt{2},$ (7)
then
$\|x+y+r\mathbf{1}_{k}\|^{2}+\|x^{\prime}+y^{\prime}+r^{\prime}\mathbf{1}_{k^{\prime}}\|^{2}\geq\varepsilon^{2}.$
To apply this lemma, first we let $k=|B|\geq 1$ and $k^{\prime}=|B^{\prime}|\geq 1$, and define
$x=w_{B}-c\mathbf{1}_{B},\ y=t^{\prime}\mathbf{1}_{B}-w^{\prime}_{B},\ x^{\prime}=w^{\prime}_{B^{\prime}}-c^{\prime}\mathbf{1}_{B^{\prime}},\ y^{\prime}=t\mathbf{1}_{B^{\prime}}-w_{B^{\prime}},$
and note that all these vectors have nonnegative entries (specifically, $x$ and $x^{\prime}$ are nonnegative by definition of $B$ and $B^{\prime}$, while $y$ and $y^{\prime}$ are nonnegative by Theorem 12, using the fact that $B\cap s(w^{\prime})=B^{\prime}\cap s(w)=\varnothing$). Then the condition (6) is satisfied by definition of $c=c_{\varepsilon}(w)$ and $c^{\prime}=c_{\varepsilon}(w^{\prime})$. Define also
$r=c-t^{\prime},\ r^{\prime}=c^{\prime}-t.$
and note that
$r+r^{\prime}=(c-t)+(c^{\prime}-t^{\prime})=\sum_{j\in[L]}(w_{j}-c)_{+}-\varepsilon/\sqrt{2}+\sum_{j\in[L]}(w^{\prime}_{j}-c^{\prime})_{+}-\varepsilon/\sqrt{2}\\ =x^{\top}\mathbf{1}_{k}+x^{\prime}{}^{\top}\mathbf{1}_{k^{\prime}}-\varepsilon\sqrt{2},$
where the second step holds by definition of $t=t_{\varepsilon}(w)$ and $t^{\prime}=t_{\varepsilon}(w^{\prime})$. Therefore, (7) is also satisfied. Returning to (5), Lemma 13 then implies that $\|w-w^{\prime}\|^{2}\geq\varepsilon^{2}$, which completes the proof of the theorem.
A.2 Proof of 9
First we verify that $\operatorname{argmax}(w)\subseteq\operatorname{argmax}^{\varepsilon}(w)$. Let $j\in\operatorname{argmax}(w)$, i.e., $w_{j}=\max_{\ell}w_{\ell}$. Let $v=w+\mathbf{e}_{j}\cdot\varepsilon/\sqrt{2}$, where $\mathbf{e}_{j}=(0,\dots,0,1,0,\dots,0)$ is the $j$th canonical basis vector. Then $v\in R_{j}^{\varepsilon}$, and $\|w-v\|=\varepsilon/\sqrt{2}<\varepsilon$, so we have $\textnormal{dist}(w,R_{j}^{\varepsilon})<\varepsilon$ and therefore $j\in\operatorname{argmax}^{\varepsilon}(w)$. We also need to verify that $\cap_{\varepsilon>0}\operatorname{argmax}^{\varepsilon}(w)=\operatorname{argmax}(w)$, i.e., for $j\not\in\operatorname{argmax}(w)$, there is some $\varepsilon>0$ with $j\not\in\operatorname{argmax}^{\varepsilon}(w)$. Let $k\in[L]$ be an index with $w_{k}>w_{j}$, and let $\varepsilon=\sqrt{2}(w_{k}-w_{j})$. Then for any $v\in R_{j}^{\varepsilon}$, we have
$\|w-v\|^{2}\geq(w_{j}-v_{j})^{2}+(w_{k}-v_{k})^{2}=(w_{k}-\varepsilon/\sqrt{2}-v_{j})^{2}+(w_{k}-v_{k})^{2}\\ \geq\inf_{t\in\mathbb{R}}\left\{\left(t-(\varepsilon/\sqrt{2}+(v_{j}-v_{k}))\right)^{2}+t^{2}\right\}=\frac{1}{2}(\varepsilon/\sqrt{2}+(v_{j}-v_{k}))^{2}\geq\varepsilon^{2},$
where the last step holds since $v_{j}-v_{k}\geq\varepsilon/\sqrt{2}$. This means that $\textnormal{dist}(w,R_{j}^{\varepsilon})\geq\varepsilon$ and so $j\not\in\operatorname{argmax}^{\varepsilon}(w)$ at this value of $\varepsilon>0$.
Next we check monotonicity in $\varepsilon$. Fix $\varepsilon<\varepsilon^{\prime}$. Then for any $j\in[L]$,
$j\in\operatorname{argmax}^{\varepsilon}(\varepsilon)\ \Longleftrightarrow\ w_{j}>t_{\varepsilon}(w)\ \Longrightarrow\ w_{j}>t_{\varepsilon^{\prime}}(w)\ \Longleftrightarrow\ j\in\operatorname{argmax}^{\varepsilon^{\prime}}(w),$
where each step holds by Theorem 12.
Next we verify monotonicity in $w$. If $w_{j}\leq w_{k}$, then by Theorem 12
$j\in\operatorname{argmax}^{\varepsilon}(w)\ \Longleftrightarrow w_{j}>t_{\varepsilon}(w)\ \Longrightarrow w_{k}>t_{\varepsilon}(w)\ \Longleftrightarrow\ k\in\operatorname{argmax}^{\varepsilon}(w).$
Finally we check permutation invariance. Suppose $v=(w_{\sigma(1)},\dots,w_{\sigma(L)})$ is a permutation of $w$. Then by construction, we have $c_{\varepsilon}(v)=c_{\varepsilon}(w)$ and $t_{\varepsilon}(v)=t_{\varepsilon}(w)$. In particular,
$j\in\operatorname{argmax}^{\varepsilon}(v)\ \Longleftrightarrow\ v_{j}>t_{\varepsilon}(v)\ \Longleftrightarrow\ w_{\sigma(j)}>t_{\varepsilon}(w)\ \Longleftrightarrow\ \sigma(j)\in\operatorname{argmax}^{\varepsilon}(w).$
A.3 Proof of 10
Define a function $f_{w}$ as
$f_{w}(c)=\left(\sum_{j\in[L]}(w_{j}-c)_{+}\right)^{2}+\sum_{j\in[L]}(w_{j}-c)^{2}_{+}.$ (8)
Then $\hat{k}(w)$ can equivalently be defined as
$\hat{k}(w)=\max\{k\in[L]:f_{w}(w_{[k]})\leq\varepsilon^{2}\}.$
(Note that this maximum is well defined, since $f_{w}(w_{[1]})=0$ and so the set is nonempty.) Since $c\mapsto f_{w}(c)$ is strictly decreasing over $c\leq w_{[1]}$, we see that the solution $c_{\varepsilon}(w)$ to the equation $f_{w}(c)=\varepsilon^{2}$ must satisfy $c_{\varepsilon}(w)\leq w_{[\hat{k}(w)]}$, and (if $\hat{k}(w)<L$) also $c_{\varepsilon}(w)>w_{[\hat{k}(w)+1]}$. In particular, this implies
$\varepsilon^{2}$ $=\left(\sum_{j\in[L]}(w_{j}-c_{\varepsilon}(w))_{+}\right)^{2}+\sum_{j\in[L]}(w_{j}-c_{\varepsilon}(w))_{+}^{2}$
$=\left(\sum_{j=1}^{\hat{k}(w)}(w_{j}-c_{\varepsilon}(w))\right)^{2}+\sum_{j=1}^{\hat{k}(w)}(w_{j}-c_{\varepsilon}(w))^{2}$
$=\left(\hat{k}(w)\cdot(\hat{A}_{1}(w)-c_{\varepsilon}(w))\right)^{2}+\left(\hat{k}(w)\hat{A}_{2}(w)-2\hat{k}(w)\hat{A}_{1}(w)c_{\varepsilon}(w)+\hat{k}(w)c_{\varepsilon}(w)^{2}\right)$
$=\hat{k}(w)(\hat{k}(w)+1)\left(c_{\varepsilon}(w)^{2}-2c_{\varepsilon}(w)\hat{A}_{1}(w)+\frac{\hat{k}(w)(\hat{A}_{1}(w))^{2}+\hat{A}_{2}(w)}{\hat{k}(w)+1}\right).$
This is a quadratic function of $c_{\varepsilon}(w)$, and is solved by
$c_{\varepsilon}(w)=\hat{A}_{1}-\sqrt{\frac{(\hat{A}_{1}(w))^{2}-\hat{A}_{2}(w)+\frac{\varepsilon^{2}}{\hat{k}(w)}}{\hat{k}(w)+1}}.$
We also have
$t_{\varepsilon}(w)$ $=c_{\varepsilon}(w)+\varepsilon/\sqrt{2}-\sum_{j\in[L]}(w_{j}-c_{\varepsilon}(w))_{+}$
$=c_{\varepsilon}(w)+\varepsilon/\sqrt{2}-\sum_{j=1}^{\hat{k}(w)}(w_{[j]}-c_{\varepsilon}(w))$
$=(\hat{k}(w)+1)c_{\varepsilon}(w)+\frac{\varepsilon}{\sqrt{2}}-\hat{k}(w)\hat{A}_{1}(w).$
Plugging in our above expression for $c_{\varepsilon}(w)$, then,
$t_{\varepsilon}(w)=\frac{\varepsilon}{\sqrt{2}}+\hat{A}_{1}(w)-\sqrt{\hat{k}(w)+1}\sqrt{(\hat{A}_{1}(w))^{2}-\hat{A}_{2}(w)+\frac{\varepsilon^{2}}{\hat{k}(w)}}.$
Since $\operatorname{argmax}^{\varepsilon}(w)=\{j:w_{j}>t_{\varepsilon}(w)\}$ by Theorem 12, this completes the proof.
A.4 Proof of 11
First we will need a lemma:
Lemma 14.
Fix $w\in\mathbb{R}^{L}$. Then $w\in R_{j}^{\varepsilon}$ if and only if $\operatorname{argmax}^{\varepsilon}(w)=\{j\}$.
For intuition on this result, we can look back at Figure 1—for instance, the blue region marked by $\{1\}$ illustrates the set of vectors $w\in\Delta_{L-1}$ such that $\operatorname{argmax}^{\varepsilon}(w)=\{1\}$, and according to this lemma, this is equal to the region $R_{1}^{\varepsilon}$.
Using this lemma, we now need to show that, for any $\varepsilon$-compatible and permutation invariant selection rule $s$, if $s(w)=\{j\}$ for some $w\in\Delta_{L-1}$ and $j\in[L]$, then $w\in R_{j}^{\varepsilon}$ for some $j$. First, we must have $j=\operatorname{argmax}(w)$, since $s$ is assumed to contain the argmax. Fix any $k\neq j$, and define a vector $v\in\Delta_{L-1}$ by permuting $w_{j}$ and $w_{k}$:
$v_{\ell}=\begin{cases}w_{k},&\ell=j,\\ w_{j},&\ell=k,\\ w_{\ell},&\ell\neq j,k.\end{cases}.$
Then by permutation invariance we have $s(v)=\{k\}$. Since $s$ is $\varepsilon$-compatible, we therefore have
$\varepsilon\leq\|w-v\|=\sqrt{(w_{j}-v_{j})^{2}+(w_{k}-v_{k})^{2}+\sum_{\ell\neq j,k}(w_{\ell}-v_{\ell})^{2}}=\sqrt{2(w_{j}-w_{k})^{2}}.$
Therefore we have $w_{j}-w_{k}\geq\varepsilon/\sqrt{2}$ for all $k\neq j$, which proves $w\in R_{j}^{\varepsilon}$ and therefore $\operatorname{argmax}^{\varepsilon}(w)=\{j\}$.
A.5 Proof of Theorem 12
Step 1: verifying that $c_{\varepsilon}(w)$ is well defined.
First we check that the solution $c_{\varepsilon}(w)$ exists and is unique. Fix $w$ and define the function $f_{w}$ as in (8). Note that $f_{w}(c)\equiv 0$ for $c\geq\max_{i}w_{i}$, and $c\mapsto f_{w}(c)$ is strictly decreasing over $c\leq\max_{i}w_{i}$, with $\lim_{c\rightarrow-\infty}f_{w}(c)=\infty$; therefore, a unique solution $c_{\varepsilon}(w)$ to the equation $f_{w}(c)=\varepsilon^{2}$ must exist.
Step 2: verifying $t_{\varepsilon}(w)\leq c_{\varepsilon}(w)$.
Next we check that the claim $t_{\varepsilon}(w)\leq c_{\varepsilon}(w)$ must hold. First, by definition of $c_{\varepsilon}(w)$, we have
$\varepsilon^{2}=\left(\sum_{j\in[L]}(w_{j}-c_{\varepsilon}(w))_{+}\right)^{2}+\sum_{j\in[L]}(w_{j}-c_{\varepsilon}(w))^{2}_{+}\leq 2\left(\sum_{j\in[L]}(w_{j}-c_{\varepsilon}(w))_{+}\right)^{2},$
where the inequality holds by the properties of the $\ell_{1}$ and $\ell_{2}$ norms. Therefore,
$\sum_{j\in[L]}(w_{j}-c)_{+}\geq\varepsilon/\sqrt{2},$
which verifies that
$t_{\varepsilon}(w)=c_{\varepsilon}(w)+\varepsilon/\sqrt{2}-\sum_{j\in[L]}(w_{j}-c_{\varepsilon}(w))_{+}\leq c_{\varepsilon}(w).$
Step 3: checking monotonicity.
Now we turn to verifying the monotonicity properties of $t_{\varepsilon}(w)$ and $c_{\varepsilon}(w)$. First we check that these functions are nonincreasing in $\varepsilon$. First, since $f_{w}(c)$ is nonincreasing in $c$, and $c_{\varepsilon}(w)$ is the solution to $f_{w}(c)=\varepsilon^{2}$, this immediately implies that $c_{\varepsilon}(w)$ is nonincreasing in $\varepsilon$. Next we turn to $t_{\varepsilon}(w)$. Define
$t^{\prime}(c)=c+\frac{1}{\sqrt{2}}\sqrt{\left(\sum_{j\in[L]}(w_{j}-c)_{+}\right)^{2}+\sum_{j\in[L]}(w_{j}-c)_{+}^{2}}-\sum_{j\in[L]}(w_{j}-c)_{+},$
so that we have $t_{\varepsilon}(w)=t^{\prime}(c_{\varepsilon}(w))$ by construction. We can verify that $c\mapsto t^{\prime}(c)$ is nondecreasing, and therefore, $t_{\varepsilon}(w)=t^{\prime}(c_{\varepsilon}(w))\leq t^{\prime}(c_{\varepsilon^{\prime}}(w))=t_{\varepsilon^{\prime}}(w)$, where the inequality holds since $c_{\varepsilon}(w)\leq c_{\varepsilon^{\prime}}(w)$.
Next we check that $t_{\varepsilon}(w)$ and $c_{\varepsilon}(w)$ are nondecreasing functions of $w$. Fix any $w\leq w^{\prime}$ (where the inequality is coordinatewise, i.e., $w_{j}\leq w^{\prime}_{j}$ for all $j\in[L]$). Since $w\mapsto f_{w}(c)$ is a nondecreasing function, we have $f_{w}(c)\leq f_{w^{\prime}}(c)$ for all $c$. We therefore have $f_{w^{\prime}}(c_{\varepsilon}(w))\geq f_{w}(c_{\varepsilon}(w))=\varepsilon^{2}=f_{w^{\prime}}(c_{\varepsilon}(w^{\prime}))$. Since $c\mapsto f_{w^{\prime}}(c)$ is nonincreasing, therefore, $c_{\varepsilon}(w^{\prime})\geq c_{\varepsilon}(w)$. Next we consider $t_{\varepsilon}$. Let $t^{\prime\prime}_{\varepsilon}(c)=c+\varepsilon/\sqrt{2}-\sum_{j\in[L]}(w_{j}-c)_{+}$, which is a nondecreasing function of $c$. Then we have $t_{\varepsilon}(w)=t^{\prime\prime}_{\varepsilon}(c_{\varepsilon}(w))\leq t^{\prime\prime}_{\varepsilon}(c_{\varepsilon}(w^{\prime}))=t_{\varepsilon}(w^{\prime})$.
Step 4: returning to the inflated argmax.
Finally, fixing any $w\in\mathbb{R}^{L}$, we will prove that $\operatorname{argmax}^{\varepsilon}(w)=\{j:w_{j}>t_{\varepsilon}(w)\}$. First choose any $j\in[L]$ with $w_{j}>t_{\varepsilon}(w)$. We need to verify that $j\in\operatorname{argmax}^{\varepsilon}(w)$. Define $v\in\mathbb{R}^{L}$ with entries
$v_{k}=\begin{cases}\max\{c_{\varepsilon}(w)+\varepsilon/\sqrt{2},w_{j}\},&k=j,\\ \min\{c_{\varepsilon}(w),w_{k}\},&k\neq j.\end{cases}$
By construction, we have $v\in R_{j}^{\varepsilon}$. We calculate
$\textnormal{dist}(w,R_{j}^{\varepsilon})^{2}$ $\leq\|w-v\|^{2}$
$=\left(w_{j}-\max\{c_{\varepsilon}(w)+\varepsilon/\sqrt{2},w_{j}\}\right)^{2}+\sum_{k\neq j}\left(w_{k}-\min\{c_{\varepsilon}(w),w_{k}\}\right)^{2}$
$=\left(c_{\varepsilon}(w)+\varepsilon/\sqrt{2}-w_{j}\right)^{2}_{+}+\sum_{k\neq j}(w_{k}-c_{\varepsilon}(w))^{2}_{+}$
$<\left(c_{\varepsilon}(w)+\varepsilon/\sqrt{2}-t_{\varepsilon}(w)\right)^{2}+\sum_{k\neq j}(w_{k}-c_{\varepsilon}(w))^{2}_{+}$
$\leq\left(\sum_{k\in[L]}(w_{k}-c_{\varepsilon}(w))_{+}\right)^{2}+\sum_{k\in[L]}(w_{k}-c_{\varepsilon}(w))^{2}_{+}=\varepsilon^{2},$
where the last two steps hold by definition of $t_{\varepsilon}(w)$ and $c_{\varepsilon}(w)$, while the strict inequality holds since $t_{\varepsilon}(w)\leq c_{\varepsilon}(w)<c_{\varepsilon}(w)+\varepsilon/\sqrt{2}$ as established above, while $w_{j}>t_{\varepsilon}(w)$ by assumption. Therefore $j\in\operatorname{argmax}^{\varepsilon}(w)$.
Now we check the converse. For this last step, we will need a lemma. The following result characterizes the projection of $w$ to the set $R_{j}^{\varepsilon}$.
Lemma 15.
Fix any $w\in\mathbb{R}^{L}$ and any $j\in[L]$. Then there is a unique $a\in\mathbb{R}$ satisfying
$w_{j}=a+\varepsilon/\sqrt{2}-\sum_{k\neq j}(w_{k}-a)_{+}.$ (9)
Moreover, defining the vector $v\in\mathbb{R}^{L}$ as
$v_{k}=\begin{cases}a+\varepsilon/\sqrt{2},&k=j,\\ a\wedge w_{k},&k\neq j,\end{cases}$ (10)
it holds that $v=\operatorname{argmin}_{u\in R_{j}^{\varepsilon}}\|w-u\|$, i.e., $v$ is the projection of $w$ to the set $R_{j}^{\varepsilon}$.
Next suppose $w_{j}\leq t_{\varepsilon}(w)$. We need to verify that $j\not\in\operatorname{argmax}^{\varepsilon}(w)$. Let $a$ and $v$ be defined as in Lemma 15 above. We can compare the equation (9) to the definition of $t_{\varepsilon}(w)$,
$t_{\varepsilon}(w)=c_{\varepsilon}(w)+\varepsilon/\sqrt{2}-\sum_{k\in[L]}(w_{k}-c_{\varepsilon}(w))_{+}=c_{\varepsilon}(w)+\varepsilon/\sqrt{2}-\sum_{k\neq j}(w_{k}-c_{\varepsilon}(w))_{+},$
where the last step holds since $w_{j}\leq t_{\varepsilon}(w)$ by assumption, and $t_{\varepsilon}(w)\leq c_{\varepsilon}(w)$ as proved above. Since $c\mapsto c+\varepsilon/\sqrt{2}-\sum_{k\neq j}(w_{k}-c)_{+}$ is an increasing function, and $w_{j}\leq t_{\varepsilon}(w)$, this implies $a\leq c_{\varepsilon}(w)$. We then calculate
$\textnormal{dist}(w,R_{j}^{\varepsilon})^{2}$ $=\|w-v\|^{2}$
[…]
$=\left(w_{j}-\left(a+\varepsilon/\sqrt{2}\right)\right)^{2}+\sum_{k\neq j}(w_{k}-a)^{2}_{+}$
[…]
$\geq\left(\sum_{k\neq j}(w_{k}-c_{\varepsilon}(w))_{+}\right)^{2}+\sum_{k\neq j}(w_{k}-c_{\varepsilon}(w))^{2}_{+}\textnormal{ since $a\leq c_{\varepsilon}(w)$}$
$=\left(\sum_{k\in[L]}(w_{k}-c_{\varepsilon}(w))_{+}\right)^{2}+\sum_{k\in[L]}(w_{k}-c_{\varepsilon}(w))^{2}_{+}\textnormal{ since $w_{j}\leq t_{\varepsilon}(w)\leq c_{\varepsilon}(w)$}$
$=\varepsilon^{2}\textnormal{ by definition of $c_{\varepsilon}(w)$.}$
Therefore, $\textnormal{dist}(w,R_{j}^{\varepsilon})\geq\varepsilon$, and so $j\not\in\operatorname{argmax}^{\varepsilon}(w)$.
A.6 Proofs of technical lemmas
A.6.1 Proof of Lemma 13
First, since $y$ is constrained to have nonnegative entries,
$\|x+y+r\mathbf{1}_{k}\|^{2}\geq\|(x+r\mathbf{1}_{k})_{+}\|^{2},$
where for a vector $v=(v_{1},\dots,v_{L})\in\mathbb{R}^{L}$, we write $(v)_{+}$ to denote the vector with $j$th entry given by $(v_{j})_{+}=\max\{v_{j},0\}$ for each $j$. The analogous bound holds for $\|x^{\prime}+y^{\prime}+r^{\prime}\mathbf{1}_{k^{\prime}}\|^{2}$.
$\|x+y+r\mathbf{1}_{k}\|^{2}+\|x^{\prime}+y^{\prime}+r^{\prime}\mathbf{1}_{k^{\prime}}\|^{2}\geq\|(x+r\mathbf{1}_{k})_{+}\|^{2}+\|(x^{\prime}+r^{\prime}\mathbf{1}_{k^{\prime}})_{+}\|^{2}.$
We now need to show that
$\|(x+r\mathbf{1}_{k})_{+}\|^{2}+\|(x^{\prime}+r^{\prime}\mathbf{1}_{k^{\prime}})_{+}\|^{2}\geq\varepsilon^{2}.$
Next let
$\bar{r}=\frac{r+r^{\prime}}{2}=\frac{x^{\top}\mathbf{1}_{k}+x^{\prime}{}^{\top}\mathbf{1}_{k^{\prime}}-\varepsilon\sqrt{2}}{2},\quad\Delta=\frac{-r+r^{\prime}}{2},$
so that we can write
$r=\bar{r}-\Delta,\ r^{\prime}=\bar{r}+\Delta$
for some $\Delta\in\mathbb{R}$. We we therefore need to show $\inf_{\Delta\in\mathbb{R}}f(\Delta)\geq\varepsilon^{2}$, where
$f(\Delta)=\|(x+(\bar{r}-\Delta)\mathbf{1}_{k})_{+}\|^{2}+\|(x^{\prime}+(\bar{r}+\Delta)\mathbf{1}_{k^{\prime}})_{+}\|^{2}.$
First, we observe that we can restrict the range of $\Delta$. Specifically, for any $\Delta\geq\max_{j\in[L]}x_{j}+\bar{r}$, we have $f(\Delta)=\|(x^{\prime}+(\bar{r}+\Delta)\mathbf{1}_{k^{\prime}})_{+}\|^{2}$, which is a nondecreasing function; therefore,
$\inf_{\Delta\geq\max_{j\in[L]}x_{j}+\bar{r}}f(\Delta)=f\left(\max_{j\in[L]}x_{j}+\bar{r}\right),$
meaning that we do not need to consider values of $\Delta$ beyond this upper bound. Applying a similar argument for a lower bound, we see that from this point on we only need to verify that
$f(\Delta)\geq\varepsilon^{2}\textnormal{ \ for $-\max_{j\in[k^{\prime}]}x^{\prime}_{j}-\bar{r}\leq\Delta\leq\max_{j\in[L]}x_{j}+\bar{r}$}.$
Moreover, for any $\Delta$, we have
$(x_{j}+(\bar{r}-\Delta))^{2}=(x_{j}+(\bar{r}-\Delta))^{2}_{+}+(x_{j}+(\bar{r}-\Delta))^{2}_{-}\leq(x_{j}+(\bar{r}-\Delta))^{2}_{+}+(\bar{r}-\Delta)^{2}$
for all $j$, by nonnegativity of $x$. Furthermore we must have
$(x_{j}+(\bar{r}-\Delta))^{2}=(x_{j}+(\bar{r}-\Delta))^{2}_{+}$
for at least one $j\in[L]$ when $\Delta\leq\max_{j\in[L]}x_{j}+\bar{r}$ as specified above (i.e., because for $j$ maximizing the entry $x_{j}$, the value $(x+(\bar{r}-\Delta)\mathbf{1}_{k})_{j}$ is nonnegative). Therefore,
$\|(x+(\bar{r}-\Delta)\mathbf{1}_{k})_{+}\|^{2}$ $\geq\|x+(\bar{r}-\Delta)\mathbf{1}_{k}\|^{2}-(k-1)(\bar{r}-\Delta)^{2}.$
Similarly, we can show that
$\|(x^{\prime}+(\bar{r}+\Delta)\mathbf{1}_{k^{\prime}})_{+}\|^{2}\geq\|x^{\prime}+(\bar{r}+\Delta)\mathbf{1}_{k^{\prime}}\|^{2}-(k^{\prime}-1)(\bar{r}+\Delta)^{2}.$
We then calculate
$f(\Delta)$ $=\|(x+(\bar{r}-\Delta)\mathbf{1}_{k})_{+}\|^{2}+\|(x^{\prime}+(\bar{r}+\Delta)\mathbf{1}_{k^{\prime}})_{+}\|^{2}$
$\geq\|x+(\bar{r}-\Delta)\mathbf{1}_{k}\|^{2}-(k-1)(\bar{r}-\Delta)^{2}+\|x^{\prime}+(\bar{r}+\Delta)\mathbf{1}_{k^{\prime}}\|^{2}-(k^{\prime}-1)(\bar{r}+\Delta)^{2}$
$=\|x\|^{2}+2(\bar{r}-\Delta)x^{\top}\mathbf{1}_{k}+(\bar{r}-\Delta)^{2}+\|x^{\prime}\|^{2}+2(\bar{r}+\Delta)x^{\prime}{}^{\top}\mathbf{1}_{k^{\prime}}+(\bar{r}+\Delta)^{2}$
$=2\varepsilon^{2}-(x^{\top}\mathbf{1}_{k})^{2}-(x^{\prime}{}^{\top}\mathbf{1}_{k^{\prime}})^{2}+2(\bar{r}-\Delta)x^{\top}\mathbf{1}_{k}+2(\bar{r}+\Delta)x^{\prime}{}^{\top}\mathbf{1}_{k^{\prime}}+2\bar{r}^{2}+2\Delta^{2},$
where the last step holds by (6).
Writing $z=x^{\top}\mathbf{1}_{k}$ and $z^{\prime}=x^{\prime}{}^{\top}\mathbf{1}_{k^{\prime}}$ for convenience, we have
$f(\Delta)$ $\geq 2\varepsilon^{2}-z^{2}-z^{\prime}{}^{2}+2(\bar{r}-\Delta)z+2(\bar{r}+\Delta)z^{\prime}+2\bar{r}^{2}+2\Delta^{2}$
$=2\varepsilon^{2}-z^{2}-z^{\prime}{}^{2}+2\left(\frac{z+z^{\prime}-\varepsilon\sqrt{2}}{2}-\Delta\right)z+2\left(\frac{z+z^{\prime}-\varepsilon\sqrt{2}}{2}+\Delta\right)z^{\prime}$
$\hskip 72.26999pt+2\left(\frac{z+z^{\prime}-\varepsilon\sqrt{2}}{2}\right)^{2}+2\Delta^{2}$
$=\varepsilon^{2}+4\left(z-\varepsilon/\sqrt{2}\right)\left(z^{\prime}-\varepsilon/\sqrt{2}\right)+2\left(\Delta-\frac{z-z^{\prime}}{2}\right)^{2}$
$\geq\varepsilon^{2}+4\left(z-\varepsilon/\sqrt{2}\right)\left(z^{\prime}-\varepsilon/\sqrt{2}\right).$
where the second step holds because $\bar{r}=\frac{r+r^{\prime}}{2}=\frac{z+z^{\prime}-\varepsilon\sqrt{2}}{2}$ by (7).
Finally, we also have $x^{\top}\mathbf{1}_{k}=\|x\|_{1}\geq\|x\|$, where the first step holds since $x$ is nonnegative, and so by assumption (6) we have $z=x^{\top}\mathbf{1}_{k}\geq\varepsilon/\sqrt{2}$. Similarly, $z^{\prime}\geq\varepsilon/\sqrt{2}$. Therefore, $4\left(z-\varepsilon/\sqrt{2}\right)\left(z^{\prime}-\varepsilon/\sqrt{2}\right)\geq 0$, and so we have shown that $f(\Delta)\geq\varepsilon^{2}$, as desired.
A.6.2 Proof of Lemma 14
First, suppose $w\in R_{j}^{\varepsilon}$. Then $\operatorname{argmax}(w)=\{j\}$, and so $j\in\operatorname{argmax}^{\varepsilon}(w)$ by 9. Next, let $c=w_{j}-\varepsilon/\sqrt{2}$. Then $w_{\ell}\leq c$ for all $\ell\neq j$, and so defining $f_{w}$ as in (8), we have
$f_{w}(c)=(w_{j}-c)^{2}+(w_{j}-c)^{2}=\varepsilon^{2}\ \Longrightarrow\ c_{\varepsilon}(w)=c=w_{j}-\varepsilon/\sqrt{2}.$
Then
$t_{\varepsilon}(w)=c_{\varepsilon}(w)+\varepsilon/\sqrt{2}-\sum_{\ell\in[L]}(w_{\ell}-c_{\varepsilon}(w))_{+}=\left(w_{j}-\varepsilon/\sqrt{2}\right)+\varepsilon/\sqrt{2}-\varepsilon/\sqrt{2}=w_{j}-\varepsilon/\sqrt{2},$
and so for all $\ell\neq j$, we have $w_{\ell}\leq w_{j}-\varepsilon/\sqrt{2}=t_{\varepsilon}(w)$ and thus $\ell\not\in\operatorname{argmax}^{\varepsilon}(w)$. This verifies that $\operatorname{argmax}^{\varepsilon}(w)=\{j\}$.
To prove the converse, suppose $\operatorname{argmax}^{\varepsilon}(w)=\{j\}$. By 9, we then have $\operatorname{argmax}(w)=\{j\}$. Let $k\in\operatorname{argmax}_{\ell\neq j}w_{\ell}$, then to show $w\in R_{j}^{\varepsilon}$ it suffices to show that $w_{j}\geq w_{k}+\varepsilon/\sqrt{2}$. Define $v\in\mathbb{R}^{L}$ as
$v_{\ell}=\begin{cases}w_{k},&\ell=j,\\ w_{k}+\varepsilon/\sqrt{2},&\ell=k,\\ w_{\ell},&\ell\neq j,k.\end{cases}.$
Then $v\in R_{k}^{\varepsilon}$, and so we must have $\|w-v\|\geq\varepsilon$ since $k\not\in\operatorname{argmax}^{\varepsilon}(w)$. We calculate
$\|w-v\|^{2}=(w_{j}-v_{j})^{2}+(w_{k}-v_{k})^{2}=(w_{j}-w_{k})^{2}+(\varepsilon/\sqrt{2})^{2}.$
This implies that we must have $(w_{j}-w_{k})^{2}\geq\varepsilon^{2}/2$, which completes the proof.
A.6.3 Proof of Lemma 15
First we verify existence and uniqueness of $a$. Let
$f(a)=w_{j}-a+\sum_{k\neq j}(w_{k}-a)_{+}.$
Then $f:\mathbb{R}\rightarrow\mathbb{R}$ is a continuous and strictly decreasing bijection, so $f(a)=\varepsilon/\sqrt{2}$ must have a unique solution.
Next let $a$ be this unique solution and let $v$ be defined as in (10). Now we verify that $v$ is the unique solution to the optimization problem
$v=\operatorname{argmin}_{u\in\mathbb{R}^{L}}\left\{\frac{1}{2}\|u-w\|^{2}:u_{j}\geq u_{k}+\varepsilon/\sqrt{2}\textnormal{ for all $k\neq j$}\right\},$
which defines the projection of $w$ to $R_{j}^{\varepsilon}$. By the Karush–Kuhn–Tucker (KKT) conditions for first-order optimality, it is sufficient to verify that
$v-w-\sum_{k\neq j}\lambda_{k}(\mathbf{e}_{j}-\mathbf{e}_{k})=0,$ (11)
where $\lambda_{k}\geq 0$ for all $k$, and $\lambda_{k}=0$ for any inactive constraints, i.e., any $k$ for which $v_{j}>v_{k}+\varepsilon/\sqrt{2}$. Let $\lambda_{k}=(w_{k}-a)_{+}$ for each $k\neq j$. Note that if $v_{j}>v_{k}+\varepsilon/\sqrt{2}$ (i.e., an inactive constraint), then we must have $w_{k}<a$ and so $\lambda_{k}=0$, by construction of $v$. Then, for any $\ell\neq j$,
$\left(v-w-\sum_{k\neq j}\lambda_{k}(\mathbf{e}_{j}-\mathbf{e}_{k})\right)_{\ell}=v_{\ell}-w_{\ell}+\lambda_{\ell}=a\wedge w_{\ell}-w_{\ell}+(w_{\ell}-a)_{+}=0,$
where $\mathbf{e}_{k}$ denotes the $k$th canonical basis vector in $\mathbb{R}^{L}$. This proves that the $\ell$th coordinate of the system of equations (11) holds for each $\ell\neq j$. Now we verify the $j$th coordinate. We have
$\left(v-w-\sum_{k\neq j}\lambda_{k}(\mathbf{e}_{j}-\mathbf{e}_{k})\right)_{j}=v_{j}-w_{j}-\sum_{k\neq j}\lambda_{k}\\ =a+\varepsilon/\sqrt{2}-w_{j}-\sum_{k\neq j}(w_{k}-a)_{+}=\varepsilon/\sqrt{2}-f(a)=0,$
by our choice of $a$ as the solution to $f(a)=\varepsilon/\sqrt{2}$. This verifies the KKT conditions, and thus, $v$ is the projection of $w$ to $R_{j}^{\varepsilon}$ as claimed.
Appendix B Extension to randomized algorithms
As mentioned in Section 2, in many applications it is common to use randomization in the construction of a classification procedure, in the learning algorithm $\mathcal{A}$ and/or in the selection rule $s$. In this section we formalize this more general framework, and will see how our results apply.
First, in the non-random setting, a learning algorithm $\mathcal{A}$ maps a data set $\mathcal{D}\in(\mathcal{X}\times[L])^{n}$ to a fitted probability estimate function $\hat{p}:\mathcal{X}\to\Delta_{L-1}$—we write the regression procedure as $\hat{p}=\mathcal{A}(\mathcal{D})$. In the randomized setting, we also allow for external randomness, $\hat{p}=\mathcal{A}(\mathcal{D};\xi)$, where $\xi\sim\textnormal{Uniform}[0,1]$ is a random seed (e.g., we might use $\xi$ to randomly shuffle the training data when running stochastic gradient descent).
Next, in the non-random setting, a selection rule $s$ is a map from $\Delta_{L-1}$ to subsets of $[L]$, resulting in a candidate set of labels $\hat{S}=s(\hat{p}(x))$. In the randomized setting (for instance, if $s$ is the argmax but with ties broken at random), we instead include a random seed $\zeta\sim\textnormal{Uniform}[0,1]$, and write $\hat{S}=s(\hat{p}(x);\zeta)$. Formally, then, we have $s:\Delta_{L-1}\times[0,1]\to\raisebox{1.79993pt}{\Large$\wp$}([L])$. (Note that the selection rule proposed in this work—the inflated argmax—is itself not random.)
Combining the two stages of the procedure, then, the classification algorithm is given by $\mathcal{C}=s\circ\mathcal{A}$, where given a training set $\mathcal{D}$ and a test point $x$, along with i.i.d. random seeds $\xi,\zeta\sim\textnormal{Uniform}[0,1]$, we return the candidate set of labels given by
$\mathcal{C}(\mathcal{D},x;\xi,\zeta)=s(\hat{p}(x);\zeta)\textnormal{ where }\hat{p}=\mathcal{A}(\mathcal{D};\xi).$
Of course, we can derive the non-random setting as a special case, simply by designing $\mathcal{A}$ and $s$ to not depend on the random seeds $\xi$ and $\zeta$.
Now, how do the results of this paper extend to this randomized setting? First, we need to reconsider our definition of selection stability (Definition 1). We will now define it as
$\frac{1}{n}\sum_{i=1}^{n}\Pr\left\{\hat{S}\cap\hat{S}^{\setminus i}=\varnothing\right\}\leq\delta,$
where $\hat{S}=\mathcal{C}(\mathcal{D},x;\xi,\zeta)$ and $\hat{S}^{\setminus i}=\mathcal{C}(\mathcal{D}^{\setminus i},x;\xi,\zeta)$, and where the probabilities are taken with respect to the distribution of the random seeds. We can also consider a notion of algorithmic stability (Definition 2) that allows for randomization: we say that $\mathcal{A}$ has tail stability $(\varepsilon,\delta)$ if
$\frac{1}{n}\sum_{i=1}^{n}\Pr\left\{\|\hat{p}(x)-\hat{p}^{\setminus i}(x)\|\geq\varepsilon\right\}\leq\delta,$
where $\hat{p}=\mathcal{A}(\mathcal{D};\xi)$ and $\hat{p}^{\setminus i}=\mathcal{A}(\mathcal{D}^{\setminus i};\xi)$, and where again probability is taken with respect to the distribution of $\xi$.
With these definitions in place, we observe that the result of 4 still holds in this more general setting for the selection rule $\operatorname{argmax}^{\varepsilon}$: if $\mathcal{A}$ is a randomized algorithm with tail stability $(\varepsilon,\delta)$ (under the new definition given above), then the randomized classification algorithm given by $\mathcal{C}=\operatorname{argmax}^{\varepsilon}\circ\mathcal{A}$ has selection stability $\delta$ (again, under the new definition given above), simply due to the $\varepsilon$-compatibility property of $\operatorname{argmax}^{\varepsilon}$ (Theorem 8). Soloff et al., 2024b proved Theorem 7 in this more general setting of randomized learning algorithms $\mathcal{A}$, so Theorem 5 holds for randomized $\mathcal{A}$ under this more general definition of selection stability.
Appendix C Extension to finitely many bags
In Section 3.1, we discussed how in practice, the bagged algorithm $\widetilde{\mathcal{A}}_{m}$ would be constructed by taking an empirical average over some large number $B$ of sampled bags, rather than computing the exact expectation $\mathbb{E}_{r}[\mathcal{A}(\mathcal{D}^{r})(x)]$. Now that we have allowed for randomized learning algorithms (as in Appendix B), we can formalize this setting: for a random seed $\xi\sim\textnormal{Uniform}[0,1]$, we define
$\widetilde{\mathcal{A}}_{m}(\mathcal{D},\xi)(x)=\frac{1}{B}\sum_{b=1}^{B}\mathcal{A}(\mathcal{D}^{r_{b}})(x),$
where the random draw of $B$ many bags, $r_{1},\dots,r_{B}$, is generated using the random seed $\xi$. Recall that Soloff et al., 2024b showed tail stability for the bagged version of any algorithm (as restated in Theorem 7)—their work also gives a result for the finite-$B$ case, as follows:
Theorem 16 ((Soloff et al., 2024b, )).
For any base learning algorithm $\mathcal{A}$ returning outputs in $\Delta_{L-1}$, the bagged algorithm $\widetilde{\mathcal{A}}_{m}$ (computed with a finite number of bags, $B$) has tail stability $(\varepsilon,\delta)$ for any $\varepsilon,\delta>0$ satisfying
$\delta=\varepsilon^{-2}\cdot(1-1/L)\left(\frac{1}{n-1}\cdot\frac{p_{n,m}}{1-p_{n,m}}+\frac{16e^{2}}{B}\right).$
In particular, combined with our result on the $\varepsilon$-compatibility of the inflated argmax, we have the following generalization of our main result (Theorem 5):
Theorem 17.
Fix any sample size $n$, any bag size $m$, any inflation parameter $\varepsilon>0$, and any number of bags $B\geq 1$. For any base learning algorithm $\mathcal{A}$, the classification algorithm $\mathcal{C}=\operatorname{argmax}^{\varepsilon}\circ\widetilde{A}_{m}$, obtained by combining the bagged version of $\mathcal{A}$ (with $B$ many bags) together with the inflated argmax, satisfies selection stability $\delta$ where
$\delta=\varepsilon^{-2}\cdot(1-1/L)\left(\frac{1}{n-1}\cdot\frac{p_{n,m}}{1-p_{n,m}}+\frac{16e^{2}}{B}\right).$ (12)
Comparing to our main result for the derandomized case (Theorem 5, which can be interpreted as taking $B\to\infty$), we see that this finite-$B$ result is essentially the same as Theorem 5 once $B\gg n$.
Appendix D Compare to a simpler selection rule: the fixed-margin rule
In this section, we compare to the simpler fixed-margin selection rule, which was defined earlier in (3)—for convenience, we define it again here:
$s_{\textnormal{margin}}^{\varepsilon}(w)=\left\{j\in[L]:w_{j}>\max_{\ell\in[L]}w_{\ell}-\varepsilon/\sqrt{2}\right\},$
the set of all indices $j$ that are within some margin of being the maximum.
In this section, we will see that this fixed-margin selection rule is $\varepsilon$-compatible. Since this rule clearly has the advantage of being much simpler than the inflated argmax (both in terms of its definition and interpretability, and in terms of its computation), we might ask whether this rule is perhaps preferable to the more complex inflated argmax. However, we will also see that the fixed-margin selection rule can be very inefficient compared to the inflated argmax: the inflated argmax can never return a larger set, and will often return a substantially smaller one.
D.1 Theoretical results
First, we verify that this simple rule satisfies $\varepsilon$-compatibility (Definition 3).
Proposition 18.
For any $\varepsilon>0$, the selection rule $s_{\textnormal{margin}}^{\varepsilon}$ is $\varepsilon$-compatible.
In particular, since $s_{\textnormal{margin}}^{\varepsilon}$ clearly contains the argmax (i.e., $\operatorname{argmax}(w)\subseteq s_{\textnormal{margin}}^{\varepsilon}(w)$) and satisfies permutation invariance (in the sense of 9), by 11 this immediately implies that
$s_{\textnormal{margin}}^{\varepsilon}(w)=\{j\}\ \Longrightarrow\ \operatorname{argmax}^{\varepsilon}(w)=\{j\}$
for all $w\in\mathbb{R}^{L}$, i.e., $\operatorname{argmax}^{\varepsilon}$ is at least as good as $s_{\textnormal{margin}}^{\varepsilon}$ at returning a singleton set. However, for this particular selection rule, we can state an even stronger result:
Proposition 19.
For any $w\in\mathbb{R}^{L}$ and any $\varepsilon>0$,
$s_{\textnormal{margin}}^{\varepsilon}(w)\supseteq\operatorname{argmax}^{\varepsilon}(w).$
Refer to captionRefer to caption
Figure 3: The fixed-margin selection rule (3), left, and the inflated argmax (2), right, when $L=3$.
In particular, this theoretical result ensures that the set of candidate labels $\hat{S}$ returned by the inflated argmax can never be larger than for the fixed-margin selection rule. In low dimensions, however, the improvement is small. Indeed, for $L=2$, we actually have $s_{\textnormal{margin}}^{\varepsilon}(w)=\operatorname{argmax}^{\varepsilon}(w)$ for all $w$. For $L=3$, Figure 3 shows that the inflated argmax is strictly better (i.e., the set inclusion result of 19 can, for certain values of $w$, be a strict inclusion), but the difference in this low-dimensional setting appears minor. Next, however, we will see empirically that as the dimension $L$ grows, the benefit of the inflated argmax can be substantial.
D.2 Simulation to compare the selection rules
In this section, we compare the fixed-margin rule to the inflated argmax using randomly generated probability weights. For our simulation, we consider the size of the two sets for various values of $L$. To sample the weights, we draw a standard Gaussian random vector $Z\sim\mathcal{N}(0,\mathbf{I}_{L})$ and define $w$ as the softmax of $Z$, i.e.,
$w_{j}=\frac{e^{Z_{j}}}{\sum_{\ell\in[L]}e^{Z_{\ell}}}$
for $j\in[L]$. We then compute $\operatorname{argmax}^{\varepsilon}(w)$ and $s_{\textnormal{margin}}^{\varepsilon}(w)$, where $\varepsilon=0.1$. In Figure 4, we present results comparing the average size of the sets returned by each of the two methods. In this setting, we see that the inflated argmax has substantial gains over the fixed-margin selection rule when the number of classes $L$ is large. Even when $L=25$, the ratio of the expected sizes is about $78\%$, so the inflated argmax has a nontrivial advantage in that setting. When $L=100$, the ratio of expected sizes is $48\%$, meaning the inflated argmax is on average more than twice as small.
Figure 4: Simulation to compare the selection rules $\operatorname{argmax}^{\varepsilon}$ and $s_{\textnormal{margin}}^{\varepsilon}$. The figure shows the average set size, $|\operatorname{argmax}^{\varepsilon}(w)|$ and $|s_{\textnormal{margin}}^{\varepsilon}(w)|$, avaeraged over $1,000$ random draws of $w$ (with standard error bars shown). See Section D.2 for details.
D.3 Proofs for the fixed-margin selection rule
D.3.1 Proof of 18
Let $w,v\in\mathbb{R}^{L}$ with $s_{\textnormal{margin}}^{\varepsilon}(w)\cap s_{\textnormal{margin}}^{\varepsilon}(v)=\varnothing$. Let $j\in\operatorname{argmax}_{\ell}w_{\ell}$ and $k\in\operatorname{argmax}_{\ell}v_{\ell}$. Then clearly $j\in s_{\textnormal{margin}}^{\varepsilon}(w)$, so since $s_{\textnormal{margin}}^{\varepsilon}(w)\cap s_{\textnormal{margin}}^{\varepsilon}(v)=\varnothing$, this implies $j\not\in s_{\textnormal{margin}}^{\varepsilon}(v)$. By definition, then, $v_{j}\leq\max_{\ell}v_{\ell}-\varepsilon/\sqrt{2}=v_{k}-\varepsilon/\sqrt{2}$. By an identical argument, we have $w_{k}\leq w_{j}-\varepsilon/\sqrt{2}$. Then
$\|w-v\|^{2}$ $\geq(w_{j}-v_{j})^{2}+(w_{k}-v_{k})^{2}$
$\geq(w_{j}-v_{j})^{2}_{+}+(v_{k}-w_{k})^{2}_{+}$
$\geq\left(w_{j}-(v_{k}-\varepsilon/\sqrt{2})\right)_{+}^{2}+\left(v_{k}-(w_{j}-\varepsilon/\sqrt{2})\right)_{+}^{2}$
$\geq\inf_{t\in\mathbb{R}}\left\{(\varepsilon/\sqrt{2}-t)_{+}^{2}+(\varepsilon/\sqrt{2}+t)_{+}^{2}\right\}$
$=2(\varepsilon/\sqrt{2})^{2}=\varepsilon^{2}.$
This proves $\|w-v\|\geq\varepsilon$, and therefore, we have shown that $\varepsilon$-compatibility is satisfied.
D.3.2 Proof of 19
Suppose $j\in\operatorname{argmax}^{\varepsilon}(w)$. Then $\textnormal{dist}(w,R_{j}^{\varepsilon})<\varepsilon$, so we can find some $v\in R_{j}^{\varepsilon}$ with $\|w-v\|<\varepsilon$. Then, for any $k\neq j$,
$w_{j}-w_{k}$ $=v_{j}-v_{k}+(w_{j}-v_{j})+(v_{k}-w_{k})$
$\geq\varepsilon/\sqrt{2}+(w_{j}-v_{j})+(v_{k}-w_{k})\textnormal{ since $v\in R_{j}^{\varepsilon}$}$
$\geq\varepsilon/\sqrt{2}-\big{(}|w_{j}-v_{j}|+|w_{k}-v_{k}|\big{)}$
$\geq\varepsilon/\sqrt{2}-\sqrt{2}\sqrt{(w_{j}-v_{j})^{2}+(w_{k}-v_{k})^{2}}$
$\geq\varepsilon/\sqrt{2}-\sqrt{2}\|w-v\|$
$>\varepsilon/\sqrt{2}-\sqrt{2}\cdot\varepsilon$
$=-\varepsilon/\sqrt{2}.$
Since this holds for all $k\neq j$, then,
$w_{j}>\max_{k\in[L]}w_{k}-\varepsilon/\sqrt{2},$
which proves that $j\in s_{\textnormal{margin}}^{\varepsilon}(w)$.